OpenStudy (anonymous):

which is an equation in form for the given point and slope. point:(7,4).. slope (6) there is multiple choice if you need

OpenStudy (imtiaz7):

Here we'll use point slope form that is: y-y1= m(x-x1) here x1=7 and y1=4 and m(slope)=6 by putting values we get: 6x-y-17=0

OpenStudy (anonymous):

omg thanks so much

OpenStudy (imtiaz7):

welcome :) @katielovesyou23